# Graph Condensation: A Survey

### v1

@@ 6.3 Federated Learning
Federated learning Fu et al. (2022) is a distributed model training framework that safeguards private data, involving multiple clients with exclusive data and a central server. In this setup, clients train local models with their unique data and then transmit the parameters to the server, which aggregates and redistributes the updated knowledge. A key challenge in federated learning is the non-i.i.d nature of client data, leading to biased local models and slow convergence, increasing the communication load between the server and clients. To address this, FedGKD Pan et al. (2023) integrates GC at the client level to efficiently extract the local task features, enhancing the similarity computation and ensuring better incorporation of local information. Combined with global aggregation, FedGKD significantly improves collaboration results in federated learning environments.
6.4 Inference Acceleration
@@
In practical graph systems, new present nodes will be connected with existing nodes in the original graph, then the message passing process is executed to aggregate information and generate representations for inference. However, the extensive size of large original graphs poses a considerable challenge for real-time inference Gao et al. (2023b). This difficulty primarily stems from the “neighbor explosion” problem, where the number of aggregated nodes exponentially increases with the propagation depth. To reduce the high computational demands and latency inherent in GNNs, MCond Gao et al. (2024) adopts GC and proposes to learn a mapping from original nodes to condensed nodes. This approach allows direct information propagation on the small condensed graph, resulting in a remarkable acceleration of the inference process.
6.5 Recommender Systems
@@
Recommender systems are widely used in various industries such as e-commerce and social media, aiming to suggest relevant items or content to users. In a typical recommender system, the user-item interactions can be represented as a bipartite graph. As the number of users and items grows, this graph becomes increasingly large and complex, posing significant computational challenges. To handle this issue, Distill-CF Sachdeva et al. (2022) applies GC to recommender systems. It adopts a multi-step Gumbel sampling trick to condense the original user base into a limited number of representative users, which are then iteratively updated through distilling knowledge from the original graph. By replacing the original user-item graph with the synthesized small graph for training and inference, this method leads to more scalable and efficient recommender systems.
@@ 7 Challenges and Future Directions
Although current GC methodologies have achieved great success across various applications, GC still grapples with inherent limitations and unresolved challenges. In this section, we summarize the current challenges and identify several future research directions worth exploration.
@@ 7.1 Condensation for Diverse Graphs
As an innovative approach in graph reduction, current GC methods have predominantly focused on simple attribute graphs. However, the graph data in practical applications is far more complex, showcasing a wide spectrum of characteristics including heterophilic graphs Zheng et al. (2022), heterogeneous graphs Gao et al. (2023a), digraphs Bang-Jensen and Gutin (2008), and dynamic graphs Kazemi et al. (2020), etc. Each type of graph brings forth its own unique features, necessitating a deep understanding of their structures and a requirement for diverse GC strategies. Such a comprehensive grasp is not only critical for effectively managing the increasing volumes of data but also plays a key role in widening the horizons of their applications.
@@ 7.2 Task-Agnostic Graph Condensation
The optimization strategies currently employed in GC are specifically designed for the classification task, heavily relying on class labels. However, the entanglement of the GC optimization process with task-related labels can potentially hinder performance in other downstream tasks. Therefore, there is a noteworthy potential to develop generalized GC methods for diverse downstream tasks, including link prediction Kumar et al. (2020), anomaly detection Akoglu et al. (2015), graph reconstruction Xia et al. (2021), and graph generation Zhu et al. (2022), etc. Furthermore, considering the extensive time and resources required for data annotation, the development of self-supervised GC can significantly advance and expand the adaptability of GC techniques.
@@ 7.3 Efficient Evaluation for Condensed Graphs
In the realm of GC, the primary objective is to create high-quality condensed graphs. However, a systematic and well-established methodology for assessing the condensed graph quality is notably absent in the field Zheng et al. (2023c). This gap is particularly critical given the complex optimization involved in GC. The assessment of the condensed graph quality is essential not only for monitoring the condensation progress but also for potentially streamlining the optimization through early-stop strategies. Currently, GC practices leverage the downstream GNNs performance as the indicator for assessment. However, this method adds considerable computational overhead to the GC process. Moreover, the concept of quality in condensed graphs is inherently complex, varying significantly across different applications and necessitating a multifaceted approach to evaluation. Considering these challenges, there is a pressing requirement to develop the comprehensive and efficient evaluation metrics. Such metrics would not only accelerate the condensation process but also broaden the range of GC applications.
7.4 Robust and Secure Graph Condensation
The condensation of graphs not only facilitates efficient training of GNNs but also contributes significantly to reduce the storage overhead and accelerate the data transmission. However, the robustness of highly compressed graph datasets to various security and privacy threats Sun et al. (2022) (e.g., adversarial attacks, inference breaches, and backdoor attacks) remains unknown. These threats pose significant limitations on the deployment of GC in security-sensitive sectors, including but not limited to confidential communication networks, financial systems and industrial control systems. In light of these challenges, the development of robust, privacy-preserving GC algorithms becomes imperative.
@@ 7.5 Explainable Graph Condensation
Another promising avenue for GC research is the development of explainable GC methods. Current GC techniques, while efficient in reducing graph size, often operate as a black box, offering limited insight into how condensed graphs represent the original graph data Yuan et al. (2022). An explainable condensation can illuminate the intricate patterns within the graph, thereby enhancing user trust and acceptance. This is particularly important in sectors like healthcare Wang et al. (2020) and finance Wang et al. (2021), where the interpretability and justification of decision-making processes are critical. Consequently, by elucidating the logic behind data transformations, explainable GC has the potential to open up new avenues for its application, establishing GC as a more reliable and transparent tool in data-centric technology fields.
@@ 8 Conclusion
This survey presents an up-to-date and comprehensive overview of GC and classifies the existing research into four distinct categories aligned with the proposed essential GC evaluation criteria. It further explores two critical components of GC: optimization strategies and condensed graph generation, facilitating a deep understanding of GC techniques. Additionally, it highlights the current challenges and emerging perspectives in GC, with the objective of inspiring and guiding future research in this evolving field. Overall, this survey seeks to serve as a beacon for researchers and practitioners, guiding them through the evolving landscape of GC and fostering advancements that could reshape the use of graph data in various practical applications.

### v2

@@ 7.1.3 Federated Learning
Federated learning [103] is a distributed framework that safeguards private data, involving multiple clients with exclusive data and a central server. In this setting, clients train local models with their unique data and then transmit the parameters to the server, which aggregates and redistributes the updated knowledge. A key challenge in federated learning is the non-i.i.d nature of client data, leading to biased local models and slow convergence, increasing the communication load between the server and clients. To address this, FedGKD [56] integrates GC at the client level to efficiently extract the local task features, enhancing the similarity computation and ensuring better incorporation of local information. Combined with global aggregation, FedGKD significantly improves collaboration results in federated learning environments. Besides integrating GC at the client level, FedGC [55] explores the federated graph condensation task, which aims to learn a unified condensed graph at the server level, encompassing knowledge from various clients. By matching the aggregated gradients from clients, FedGC tackles the data heterogeneity problem and generates a global condensed graph, facilitating efficient model training in the cross-silo scenarios.
7.1.4 Inference Acceleration
@@
In practical graph systems, new present nodes will be connected with existing nodes in the original graph, then the message-passing process is executed to aggregate information and generate representations for inference. However, the extensive size of large original graphs poses a considerable challenge for real-time inference [35]. This difficulty primarily stems from the “neighbor explosion” problem, where the number of aggregated nodes exponentially increases with the propagation depth. To reduce the high computational demands and latency inherent in GNNs, MCond [26] adopts GC and proposes to learn a mapping from original nodes to condensed nodes. This approach allows direct information propagation on the small condensed graph, resulting in a significant acceleration of the inference process.
7.1.5 Recommender Systems
@@
Recommender systems are widely used in various industries such as e-commerce and social media, aiming to suggest relevant items or content to users. In a typical recommender system, the user-item interactions can be represented as a bipartite graph. As the number of users and items grows, this graph becomes increasingly large and complex, posing significant computational challenges. To handle this issue, Distill-CF [104] applies GC to recommender systems. It adopts a multi-step Gumbel sampling trick to condense the original user base into a limited number of representative users, which are then iteratively updated through distilling knowledge from the original graph. By replacing the original user-item graph with the synthesized small graph for training and inference, this method leads to more scalable and efficient recommender systems.
7.1.6 Heterogeneous Graphs
Heterogeneous graphs [105, 106] are adept at modeling complex systems characterized by multi-typed entities and relations, thereby providing a broad range of practical applications and significant potential for generalization. However, their complex structure and inherent heterogeneity present significant challenges in data management and in training heterogeneous GNNs. In light of these challenges, HGCond [53] extends GCond [25] from simple graphs to heterogeneous graphs, generating a condensed heterogeneous graph for efficient model training. To address the heterogeneity problem, HGCond initializes condensed nodes by clustering same-type original nodes and further enhances the gradient matching optimization by constraining the parameter space of the relay model with orthogonal parameter sequences. This approach significantly enhances the scalability and generalization of heterogeneous GNNs on heterogeneous graphs, thereby broadening their applications in managing complex, large-scale systems.
7.2 Open-Source Libraries
Several notable PyTorch-based GC libraries have been developed based on the widely-used GNN libraries PyTorch Geometric (PyG) [^3] https://www.pyg.org/ and Deep Graph Library (DGL) [^4] https://www.dgl.ai/. These libraries provide essential tools and benchmarks to advance research and applications in graph-based systems.
  • GCondenser [^5] https://github.com/superallen13/GCondenser [44], compatible with both PyG and DGL, supports 6 GC methods evaluated across 6 diverse graph datasets for node classification tasks. It also includes implementations for applying GC methods to graph continual learning and hyperparamter sweeping by Optuna [107].
  • GC-Bench [^6] https://github.com/RingBDStack/GC-Bench [46] is a unified GC benchmark based on PyG. It incorporates 9 GC methods assessed over 12 graph datasets for node- and graph-level tasks, including node/graph classification, link prediction, node clustering, and anomaly detection.
  • GraphSlim [^7] https://github.com/Emory-Melody/GraphSlim [45] is a PyG-based library for graph reduction, including sparsification, coarsening, and condensation. It evaluates 7 GC methods on 5 graph datasets for node classification and extends the utility of GC to neural architecture search. Additionally, it assesses the robustness of GC methods by integrating the DeepRobust [108] library.
@@ 8 Challenges and Future Directions
Although current methodologies in GC have shown remarkable performance across various applications, the field still grapples with inherent limitations and unresolved challenges. In this section, we summarize the current challenges and identify several future research directions worth exploring.
@@ 8.1 Condensation for Diverse Graphs
As an innovative approach for graph reduction, current GC methods have predominantly focused on simple graphs. However, the graph data in practical applications is far more complex, showcasing a wide spectrum of characteristics including heterophilic graphs [109], digraphs [110], and dynamic graphs [111], etc. Each type of graph brings forth its own unique features, necessitating a deep understanding of their structures and a requirement for diverse GC strategies. Such a comprehensive grasp is not only critical for effectively managing the increasing volumes of data but also plays a key role in widening the horizons of their applications.
@@ 8.2 Task-Agnostic Graph Condensation
The optimization strategies currently employed in GC are specifically designed for the classification task, heavily relying on class labels. However, the entanglement of the GC optimization process with task-related labels can potentially hinder performance in other downstream tasks. Therefore, there is a noteworthy potential to develop generalized GC methods for diverse downstream tasks, including link prediction [112], anomaly detection [113], graph reconstruction [114], and graph generation [115], etc. Furthermore, considering the extensive time and resources required for data annotation, the development of self-supervised GC can significantly advance and expand the adaptability of GC techniques.
@@ 8.3 Efficient Evaluation for Condensed Graphs
In the realm of GC, the primary objective is to create high-quality condensed graphs. However, a systematic and well-established methodology for assessing the condensed graph quality is notably absent in the field [28]. This gap is particularly critical given the complex optimization involved in GC. The assessment of the condensed graph quality is essential not only for monitoring the condensation progress but also for potentially streamlining the optimization through early-stop strategies. Currently, GC practices leverage the downstream GNNs’ performance as the indicator for assessment. However, this method adds considerable computational overhead to the GC process. Moreover, the concept of quality in condensed graphs is inherently complex, varying significantly across different applications and necessitating a multifaceted approach to evaluation. Considering these challenges, there is a pressing requirement to develop comprehensive and efficient evaluation metrics. Such metrics would not only accelerate the condensation process but also broaden the range of GC applications.
8.4 Secure Graph Condensation
The condensation of graphs not only facilitates efficient training of GNNs but also contributes significantly to reducing the storage overhead and accelerating data transmission. However, the security of highly compressed graph datasets to various privacy threats [116] remains under-explored. Recent study [88] has exposed vulnerabilities of condensed graphs to backdoor attacks, while additional threats such as adversarial attacks and inference breaches present substantial challenges for deploying GC in security-sensitive sectors, including confidential communications, financial systems, and industrial controls. In light of these challenges, the development of secure, privacy-preserving GC algorithms becomes imperative.
@@ 8.5 Explainable Graph Condensation
Another promising avenue for GC research is the development of explainable GC methods. Current GC techniques, while efficient in reducing graph size, often operate as a black box, offering limited insight into how condensed graphs represent the original graph data [117]. An explainable condensation can illuminate the intricate patterns within the graph, thereby enhancing user trust and acceptance. This is particularly important in sectors like healthcare [118] and finance [33], where the interpretability and justification of decision-making processes are critical. Consequently, by elucidating the logic behind data transformations, explainable GC has the potential to open up new avenues for its application, establishing GC as a more reliable and transparent tool in data-centric technology fields.
@@ 9 Conclusion
This survey presents an up-to-date and comprehensive overview of GC and classifies the existing research into five distinct categories aligned with the proposed essential GC evaluation criteria. It further explores two critical components of GC: optimization strategies and condensed graph generation, facilitating a deep understanding of GC techniques. Additionally, it highlights the current challenges and emerging perspectives in GC, with the objective of inspiring and guiding future research in this evolving field. Overall, this survey seeks to serve as a beacon for researchers and practitioners, guiding them through the evolving landscape of GC and fostering advancements that could reshape the use of graph data in various practical applications.